Fermi National Laboratory


Accelerator Update

Monday January 19
The day shift began with Operations monitoring store 3179, with the Antiproton Source (Pbar) stacking, with the Recycler holding on to a stash of antiprotons, and with MiniBooNE taking beam.

Operators noted that the chilled water pressures for Pbar rose significantly at 9:34 AM. They notified the Duty Mechanic who had the pressures back to normal by 10:12 AM.

At 9:41 AM, a Switchyard expert began a study period.

Operations reported at 11:21 AM that since returning to stacking on the midnight shift, Pbar continued to have problems with its stacktail cooling system. An expert was working on the problem.

At 11:25 AM, Operations reported that the I- Source was continuing to spark on occasion.

MiniBooNE had tripped off four times by 12:46 PM due to a wandering power supply.

Switchyard experts completed their studies at 4:46 PM.

Operations noted some problems with the beam for the Collider Detector Facility (CDF) at 10 PM. They believed it was due to a collimator in Tevatron (TeV) sector E0 that had not moved to its proper position. Operators called an expert for advice.

At 10:19 PM, a TeV RF station (TRF2) tripped off due to a water skid failure. Operators consulted with experts and were advised to do nothing until after the termination of the store. Anything they do might cause the beam to abort.

Tuesday January 20
The midnight shift began with Operations monitoring store 3179, with Pbar stacking, and with MiniBooNE taking beam.

At 12:47 AM, Operations successfully moved the TeV E0 collimator into position.

Operations terminated store 3179 at 8 AM.

At 10:33 AM, technicians reported changing a pump motor on TRF2 and a power amplifier on TRF7. However, there were still problems with the ability to turn gates on and off.

Operations put the TeV through a dry squeeze at 11:47 AM.

At 2:04 PM, Operations had put the TeV into shot setup and was in the process of scraping the beam when the TeV suffered a sector F4 quench. Pbar resumed stacking.

A Linac RF station (KRF5) tripped off due to a spark indication at 2:21 PM. After a number of resets operators contacted an expert who got the station on. Then at 2:38 PM it tripped off again and would not reset due to a phase problem. This problem resolved itself.

At 3:35 PM, operators made CDF and D-zero (D0) ready for an hour access.

Cryo system techs at CHL reported a major power glitch at 4:54 PM, which tripped off their helium system. They were offline, but power had returned. The Wideband Lab reported seeing a power glitch also. Operators contacted the high voltage expert.

At 5:24 PM, the Duty Mechanic reported that feeder 30 had tripped.

Operations reported at 5:49 PM that earlier on shift Booster personnel changed out a power supply that was causing beam movement in the Main Injector (MI) and MiniBooNE.

At 7:48 PM, the high voltage expert reported that the feeder fault was in an oil switch that fed part of the Meson area. He didn't believe that this problem could be quickly resolved.

The high voltage expert bypassed the problem buildings in Meson and then energized feeder 30. Operators informed the building managers of those areas without power.

Wednesday January 21
The midnight shift began with the TeV recovering from a quench, with Pbar stacking, and with MiniBooNE taking beam.

At 12:30 AM, CHL reported seeing another power glitch that tripped off their nitrogen system this time. The Duty Electrician investigated the building that lost power. At 1:19 AM, he said he wouldn't reset feeder 30.

Operations established store 3183 at 3 AM. It's initial luminosity was 36.9E30. Pbar resumed stacking.

By 6:45 AM, Recycler had received shots of antiprotons from Pbar. These transfers were about 88% efficient.

Today's Plans
The plan for today is to switch the CHL and Meson power to a different transformer.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/21/2004   email Fermilab